Please help me write 3(cos 60 degrees + i sin 60 degrees) in the form of a + bi!
Thank you, Tami

Answer by Alan3354(69443)

Please help me write 3(cos 60 degrees + i sin 60 degrees) in the form of a + bi!
Thank you, Tami
--------------------------
3(cos 60 degrees + i sin 60 degrees)
cos(60º) = 1/2
sin(60º) = (1/2)sqrt(3)
= 3[(1/2) + i*sqrt(3)/2]
= 3/2 + i*(3*sqrt(3)/2)
a = 3/2
b = (3/2)*sqrt(3)